OpenStudy (anonymous):

simplify by factoring square root 150

OpenStudy (anonymous):

It wants you to break up 150 into a product of prime numbers. For example, 25 = 5 * 5, and so you know that sqrt(25) = 5. And 75 = 3 * 25 = 3 * 5 * 5. So sqrt(75) = sqrt(3) * 5.

OpenStudy (anonymous):

Do something similar for 150. First find what product of prime numbers gives 150.

OpenStudy (anonymous):

150=25*6 i think u know what is next

OpenStudy (anonymous):

so it would be 5 sqrt6

OpenStudy (anonymous):

???

OpenStudy (anonymous):

since u can not simplify it anymore that's ur answer

OpenStudy (anonymous):

well done

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

it's my pleasure